OpenStudy (anonymous):

Evaluate f(x) = 8 sin x − 6 cos 2x at x = π/6. 1. f(π/6) = 3√3 2. f(π/6) = 1 3. f(π/6) = 2√3 4. f(π/6) = 2 5. f(π/6) = √3 6. f(π/6) = 3

OpenStudy (anonymous):

\[\sin \left( \frac{ \pi }{6 } \right) = \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

\[\cos \left( \frac{ 2\pi }{ 6} \right) = \cos \left( \frac{ \pi }{ 3} \right) = \frac{ 1 }{ 2}\]

OpenStudy (anonymous):

We know this from the unit circle.

OpenStudy (anonymous):

\[8\left( \frac{ 1 }{ 2 } \right) - 6\left( \frac{ 1 }{ 2 } \right)\]

OpenStudy (anonymous):

4-3 = 1